OpenStudy (anonymous):

(-5,-5sqr root 3) to polar coordinates

OpenStudy (raden):

r = sqrt(x^2+y^2) tan(alpha) = y/x solve for alpha

OpenStudy (raden):

here given x=-5, y=-5sqrt(3)